OpenStudy (anonymous):

I'm really bad at fractions Solve for x: 2 over 5 (x - 2) = 4x. 2 over 9 9 negative 2 over 9 negative 9 over 2

OpenStudy (anonymous):

@lacrosseplayer22

OpenStudy (anonymous):

\[\frac{ 2 }{ 5 }(x-2)=4x\]

OpenStudy (anonymous):

thxs for typeing the equation

OpenStudy (anonymous):

\[\frac{ 2 }{ 5x-2 }\] ? or \[\frac{ 2 }{ 5 }(x-2)\]

OpenStudy (anonymous):

kio was right

OpenStudy (anonymous):

Multiple both sides by 5/2 then continue to simplify.

OpenStudy (anonymous):

So you will have\[x-2 = \frac{ 20x }{ 2 }\]

OpenStudy (anonymous):

ok i kinda get it

OpenStudy (anonymous):

2/5(x-2)=4x 2/5x-10=4x 2(5x-10)/5x-10=4x(5x-10)= 2=20x^2-40x 20x^2-40x-2=0 then you factor that

OpenStudy (anonymous):

never mind you can do the quadratic formula don't factor cause I don't think its possible

OpenStudy (anonymous):

lol i was likw wut?

OpenStudy (anonymous):

x-2 = 10x => -9x = 2 => x = -2/9

OpenStudy (anonymous):

so c?

OpenStudy (anonymous):

@lacrosseplayer22 You should have it as 2/5x - 4/5 = 4x when you distribute.

OpenStudy (anonymous):

@whereismymind49 wouldnt it be positive because both 9 and 2 and negative

OpenStudy (anonymous):

No when you add over the two

OpenStudy (anonymous):

nevermind didnt see the 20/x was looking at -9